### v1

@@ 4 Proof of Constant Competitive Ratio
First, let $\{\mathcal{X}_{k}\}$ denote the sequence of batches iteratively generated in Phase 1 of Sorted-F. To construct $\textup{{Sorted-F}}_{\mathrm{separate}}$, we process $\{\mathcal{X}_{k}\}$ one at a time. Specifically, for any batch $\mathcal{X}_{k}$, all requests in it start simultaneously, and the subsequent batch $\mathcal{X}_{k+1}$ does not begin until all requests in $\mathcal{X}_{k}$ completed. This sequential batch processing simulates a worst-case scenario by ensuring that no request from subsequent batches is initiated until the preceding batch has fully finished.
@@
For any list of requests $\mathcal{I}$, the total end-to-end latency of Sorted-F is bounded above by that of $\text{{Sorted-F}}_{\text{separate}}$, i.e.,
@@
For any list of requests $\mathcal{I}$, there exists a grouping strategy $\text{{Sorted-F}}_{\text{group}}$ that aggregates $\{\mathcal{X}_{k}\}$ into larger groups $\{\mathcal{Y}_{m}\}$ such that the terminal batch of each group nearly saturates memory $M$. Additionally, the total end-to-end latency of $\text{{Sorted-F}}_{\text{separate}}$ is bounded above by
@@
Therefore, given any list of requests $\mathcal{I}$,
@@
For any list of requests $\mathcal{I}$, the total end-to-end latency of $\text{{Sorted-F}}_{\text{group}}$ is bounded above by
@@
Subsequently, given any list of requests $\mathcal{I}$,
@@
For any list of requests $\mathcal{I}$, there exists a transformed schedule $\text{{Optimal}}_{\text{align}}$ with time-window scaling and response-length aligning that satisfies
@@
where $d_{g}=\sum_{j=1}^{g}c_{j}$ counts the cumulative batches up to group $g$, with $c_{j}$ denoting the number of batches in $\mathcal{Z}_{j}$. The decomposition proceeds by first sorting all requests in $\mathcal{Z}_{g}$ by their original initiation times. Then, for each subsequent batch $\mathcal{W}_{d_{g-1}+k}$, we select the maximal subset of remaining requests with the earliest initiation times while ensuring the total memory requirement does not exceed capacity $M$. This process repeats until all requests in $\mathcal{Z}_{g}$ are assigned to batches. This process guarantees that all batches except possibly the final one ($\mathcal{W}_{d_{g}}$) nearly saturate memory $M$ by Definition 4.4.
@@
Therefore, given any list of requests $\mathcal{I}$, we establish a lower bound on Optimal’s total end-to-end latency through the following derivation:
@@
For any list of requests $\mathcal{I}$, the total end-to-end latency of $\text{{Sorted-F}}_{\text{align}}$ is upper bounded by the total end-to-end latency of $\text{{Optimal}}_{\text{align}}$, i.e.,
@@
Proof of Theorem 4.1: Given any list of requests $\mathcal{I}$, we combine the guarantees from Theorems 4.2, 4.5, 4.8, 4.10, and 4.12 to derive the following chain of inequalities for the total end-to-end latency:
$\text{TEL}(\text{{Sorted-F}};\mathcal{I})$ $\leq\text{TEL}(\text{{Sorted-F}}_{\text{separate}};\mathcal{I})<4\cdot\text{TEL}(\text{{Sorted-F}}_{\text{group}};\mathcal{I})<8\cdot\text{TEL}(\text{{Sorted-F}}_{\text{align}};\mathcal{I})$
$\leq 8\cdot\text{TEL}(\text{{Optimal}}_{\text{align}};\mathcal{I})\leq 48\cdot\text{TEL}(\text{{Optimal}};\mathcal{I}),$

### v2

@@ 4 Proof of Constant Competitive Ratio
Before delving into the technical steps, we clarify the asymptotic regime in which the analysis is carried out. Throughout the proof, we consider a large-instance setting where the request set $\mathcal{I}$ is sufficiently large relative to the memory budget $M$. Recall from Section 2 that each request has size $s_{i},o_{i}=\epsilon(M)$, meaning that individual requests consume a vanishing fraction of memory. W.L.O.G, we further assume that the total number of requests $n=n(M)$ satisfies $\frac{n(M)\,\epsilon(M)}{M}\;\longrightarrow\;\infty$ as $M\to\infty$. Otherwise, by processing the finite amount of requests, any scheduling algorithm can have a constant competitive ratio. In words, the request set is effectively infinite, and its cardinality grows much faster than the maximum number of requests that can fit into a single batch of capacity $M$. This ensures that the contribution of any fixed group becomes negligible compared with the “tail” of remaining requests, thereby justifying the asymptotic bounds used in later arguments.
With this standing assumption in place, we now begin the first step of the proof. Let $\{\mathcal{X}_{k}\}$ denote the sequence of batches iteratively generated in Phase 1 of Sorted-F. To construct $\textup{{Sorted-F}}_{\mathrm{separate}}$, we process $\{\mathcal{X}_{k}\}$ one at a time. Specifically, for any batch $\mathcal{X}_{k}$, all requests in it start simultaneously, and the subsequent batch $\mathcal{X}_{k+1}$ does not begin until all requests in $\mathcal{X}_{k}$ completed. This sequential batch processing simulates a worst-case scenario by ensuring that no request from subsequent batches is initiated until the preceding batch has fully finished.
@@
For any $\mathcal{I}$, the total end-to-end latency of Sorted-F is bounded above by that of $\text{{Sorted-F}}_{\text{separate}}$, i.e.,
@@
For any $\mathcal{I}$, there exists a grouping strategy $\text{{Sorted-F}}_{\text{group}}$ that aggregates $\{\mathcal{X}_{k}\}$ into larger groups $\{\mathcal{Y}_{m}\}$ such that the terminal batch of each group nearly saturates memory $M$. Additionally, the total end-to-end latency of $\text{{Sorted-F}}_{\text{separate}}$ is bounded above by
@@
Therefore, given any $\mathcal{I}$,
@@
For any $\mathcal{I}$, the total end-to-end latency of $\text{{Sorted-F}}_{\text{group}}$ is bounded above by
@@
Subsequently, given any $\mathcal{I}$,
@@
For any $\mathcal{I}$, there exists a transformed schedule $\text{{Optimal}}_{\text{align}}$ with time-window scaling and response-length aligning that satisfies
@@
where $d_{g}=\sum_{j=1}^{g}c_{j}$ counts the cumulative batches up to group $g$, with $c_{j}$ denoting the number of batches in $\mathcal{Z}_{j}$. The decomposition proceeds by first sorting all requests in $\mathcal{Z}_{g}$ by their original initiation times. Tnhen, for each subsequent batch $\mathcal{W}_{d_{g-1}+k}$, we select the maximal subset of remaining requests with the earliest initiation times while ensuring the total memory requirement does not exceed capacity $M$. This process repeats until all requests in $\mathcal{Z}_{g}$ are assigned to batches. This process guarantees that all batches except possibly the final one ($\mathcal{W}_{d_{g}}$) nearly saturate memory $M$ by Definition 4.4.
@@
Therefore, given any $\mathcal{I}$, we establish a lower bound on Optimal’s total end-to-end latency through the following derivation:
@@
For any $\mathcal{I}$, the total end-to-end latency of $\text{{Sorted-F}}_{\text{align}}$ is upper bounded by the total end-to-end latency of $\text{{Optimal}}_{\text{align}}$, i.e.,
@@
Proof of Theorem 4.1: Given any $\mathcal{I}$, we combine the guarantees from Theorems 4.2, 4.5, 4.8, 4.10, and 4.12 to derive the following chain of inequalities for the total end-to-end latency:
$\text{TEL}(\text{{Sorted-F}};\mathcal{I})\leq\text{TEL}(\text{{Sorted-F}}_{\text{separate}};\mathcal{I})<4\cdot\text{TEL}(\text{{Sorted-F}}_{\text{group}};\mathcal{I})$
$<8\cdot\text{TEL}(\text{{Sorted-F}}_{\text{align}};\mathcal{I})\leq 8\cdot\text{TEL}(\text{{Optimal}}_{\text{align}};\mathcal{I})\leq 48\cdot\text{TEL}(\text{{Optimal}};\mathcal{I}),$